OpenStudy (anonymous):

help 6^3√(16z^7)+4√(200z^5)

jimthompson5910 (jim_thompson5910):

Is this \[\Large 6^3\sqrt{16z^7}+4\sqrt{200z^5}\] OR is it \[\Large 6\sqrt[3]{16z^7}+4\sqrt{200z^5}\]

OpenStudy (anonymous):

the second one

jimthompson5910 (jim_thompson5910):

alright thank you

OpenStudy (anonymous):

thankyou

jimthompson5910 (jim_thompson5910):

\[\Large 6\sqrt[3]{16z^7}+4\sqrt{200z^5}\] \[\Large 6\sqrt[3]{8*2z^6*z}+4\sqrt{100*2z^4*z}\] \[\Large 6\sqrt[3]{8z^6*2z}+4\sqrt{100z^4*2z}\] \[\Large 6\sqrt[3]{2^3(z^2)^3*2z}+4\sqrt{10^2(z^2)^2*2z}\] \[\Large 6\sqrt[3]{(2z^2)^3*2z}+4\sqrt{(10z^2)^2*2z}\] \[\Large 6\sqrt[3]{(2z^2)^3}*\sqrt[3]{2z}+4\sqrt{(10z^2)^2}*\sqrt{2z}\] \[\Large 6(2z^2)*\sqrt[3]{2z}+4(10z^2)*\sqrt{2z}\] \[\Large 12z^2\sqrt[3]{2z}+40z^2\sqrt{2z}\] ----------------------------------------------- So \[\Large 6\sqrt[3]{16z^7}+4\sqrt{200z^5}\] simplifies to \[\Large 12z^2\sqrt[3]{2z}+40z^2\sqrt{2z}\]

jimthompson5910 (jim_thompson5910):

unfortunately we can't go any further because we can't combine the cube and square roots, so we stop here